Minute of KAGRA commissioning meeting
- Date: 2018/10/30 11AM~12PM
- Attendance: Miyakawa, Oshino, Yokozawa
- eZuce: Enomoto, Kokeyama, Aso, Akutsu, Miyoko, Ushiba, Uchiyama, Miyoki, Aso
- 11/5-end of Nov.: IYA root will be closed.
- 11/6: AM or PM for 30min : Inspection for ETMX cleaning
- 12/10 or 11: ETMX cleaning
- MZ work needs stopping IR in the daytime; ~10/31 ok, 11/1~11/31?
- Assignment of tuning for suspensions.
- X arm lock
- OLTF was measured. Results of cavity pole and finesse are quite coincident with measumrements of mirrors.
- Bottom one is what we suspect that the magnet might be broken.
- Spacer can be seen.
- Magnet size is 2mm in diameter and 2mm in length, so it is difficult to think something hit the magnet and broken. It might be caused drastic temperature change.
- Laser, PSL, IOO report
- AOM driver was replaced and now AOM is working well. So, it won't change schedule.
- OMC installation
- They saw some resonance. Needs to fix some circuit for DCPD.
- Scheduling for maintenance time on Wednesday morning. All suspensions go to the safe mode.
- Picomotor driver seemed to be fixed. Needs final check with what had no response at all before.
- Next meeting: 11/6.